OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8) @TheSmartOne @SwaqqOut_KID1

OpenStudy (anonymous):

Another method to try is point-slope form. y-y1= m (x - x1) (x1, y1) is your point and m is your slope. So input the information you are given into the formula above. y-(2)= 4(x-(-1)) Then you simplify. y-2= 4(x+1) y-2 = 4x+4 And then you solve for y. (In other words, put the equation into slope-intercept form, y=mx+b. Your slope is m and your intercept is b) y = 4x + 6 From this we can see that our slope is 4, and our y-intercept is (0,6).

OpenStudy (anonymous):

thank you